# Causal Network Recovery in Perturb-seq Experiments Using Proxy and Instrumental Variables

### v1

@@ Abstract
Emerging single-cell technologies that integrate CRISPR-based genetic perturbations with single-cell RNA sequencing, such as Perturb-seq, have substantially advanced our understanding of gene regulation and causal influence of genes. While Perturb-seq data provide valuable causal insights into gene–gene interactions, statistical concerns remain regarding unobserved confounders that may bias inference. These latent factors may arise not only from intrinsic molecular features of regulatory elements encoded in Perturb-seq experiments, but also from unobserved genes arising from cost-constrained experimental designs. Although methods for analyzing large-scale Perturb-seq data are rapidly maturing, approaches that explicitly account for such unobserved confounders in learning the causal gene networks are still lacking. Here, we propose a novel method to recover causal gene networks from Perturb-seq experiments with robustness to arbitrarily omitted confounders. Our framework leverages proxy and instrumental variable strategies to exploit the rich information embedded in perturbations, enabling unbiased estimation of the underlying directed acyclic graph (DAG) of gene expressions. Simulation studies and analyses of CRISPR interference experiments of K562 cells demonstrate that our method outperforms baseline approaches that ignore unmeasured confounding, yielding more accurate and biologically relevant recovery of the true gene causal DAGs.
@@ 1 Introduction
A central objective in modern genomics is to recover the directed structure of gene–gene regulation that governs cellular function. Single-cell CRISPR perturbation assays coupled with scRNA-seq, popularized by Perturb-seq, deliver interventional readouts at single-cell resolution across thousands of targeted knockdowns or activations in pooled format (Dixit2016PerturbSeq, replogle2022mapping). Beyond revealing heterogeneous and context-specific responses at the single cell levels, these technologies create large-scale datasets that, in principle, permit causal discovery of causal gene networks (CGNs) rather than mere association maps. Such CGNs are often presented as directed acyclic graphs (DAGs), where nodes are genes and directed edges present regulatory relationships.
@@
From a statistical perspective, learning DAGs from observational or interventional data has been studied widely in literature. Classical constraint-based procedures such as the PC algorithm (Spirtes2000Causation) rely on conditional independence tests to identify Markov equivalence classes under assumptions of causal sufficiency and faithfulness. Score-based approaches such as the Greedy Equivalence Search (GES) (Chickering2002Optimal) optimize penalized likelihoods over equivalence classes, while hybrid approaches combine constraint and score elements. Continuous optimization frameworks such as NOTEARS (Zheng2018DAGs) and its nonlinear extensions (Zheng2020Nonlinear) recast DAG learning as smooth optimization problems, enabling scalable inference in high dimensions. Other methods, including those based on additive noise models (Peters2014Causal), provide identifiability for non-Gaussian or nonlinear settings. While these frameworks have advanced causal discovery in observational data, they generally assume that all confounders are observed and that interventions are perfectly applied and measured. When interventional data are available, causal discovery can, in principle, be strenghtened. Algorithms such as GIES (Hauser2012Characterization) extend GES to integrate intervention targets for orienting edges and improving identifiability, and active-learning variants optimize which interventions to perform next (Agrawal2019ABCD).
@@
However, these methods typically presume that all participating nodes of the underlying DAG are observed in the data, without omitted factors that can alter the distributions of confounded nodes. In pooled CRISPR single-cell screens, this assumption is likely to be violated. In particular, biological heterogeneity, including cell cycle phase, chromatin accessibility, and lineage differences, introduces additional latent variation that is correlated with the transcriptomic profiles of cells. Moreover, due to cost considerations, researchers implementing CRISPR experiments are often interested in intervening on a small set of key genes, which can produce incomplete datasets with omitted genes. Ignoring unobserved factors may result in biased inference of causal relationships in DAGs. Other challenges remain. Each cell can carry a random number of detected guide RNAs, even at low multiplicity of infection, so perturbation assignment is measured with error. Technical covariates such as sequencing depth, capture efficiency, and batch effects influence both gene expression and gRNA detection. Moreover, the data are high-dimensional and highly unbalanced across targets, with a small number of cells per perturbation and unknown timescales for transcriptional responses. These characteristics violate assumptions underlying standard DAG-learning methods.
@@
Several computational tools have been proposed to model Perturb-seq data. The original Perturb-seq studies introduced MIMOSCA for regularized linear modeling of perturbation effects (Dixit2016PerturbSeq). Subsequent methods such as scMAGeCK (Yang2019scMAGeCK) extended genotype-to-multivariate-phenotype mapping for single-cell CRISPR screens. More recently, SCEPTRE (Katsevich2022SCEPTRE) provided a statistically calibrated approach using conditional resampling to correct for confounding from technical covariates, substantially improving type I error control and sensitivity. Despite these advances, current frameworks primarily focus on estimating pairwise perturbation–response effects and do not naturally extend to reconstructing global regulatory DAGs. Moreover, robust computational methods for estimating DAGs in the presence of unobserved factors are still lacking.
@@
In this paper, we propose ARGEN for Arbitrary-confounder Robust causal GEne Network, as a statistical framework for recovering causal network among genes from Perturb-seq experiments that is robust to arbitrarily omitted confounders and explicitly leverages the interventional design of CRISPR screens. Our approach adopts an instrumental variable strategy that uses proxy gene expression to identify causal gene networks. Guide detections serve as instrumental variables, and proxies for hidden gene expression provide robustness against confounding. We extend the classical measurement model for scRNA-seq data (townes2019feature, sarkar2021separating) to a setting that explicitly reflects both the underlying gene regulatory network and the perturbation structure inherent to Perturb-seq data. Specifically, we employ a structural equation model (SEM) with embedded perturbation terms to represent the causal relationships among gene expressions under CRISPR interventions. Under the assumed perturbation model, we investigate conditions for identification of the ancestors and descendants of each perturbed (intervened) target gene, as well as the identification of the causal ordering and the underlying DAG. We further develop computationally efficient algorithms for estimating descendant sets and parental relationships, with provable guarantees that the resulting estimated structure satisfies the acyclicity constraint and yields a valid DAG, even in finite-sample settings.
@@ 2 Statistical Model for Perturb-seq Data
Figure 1: Schematic overview of ARGEN. a. ARGEN takes as input single-cell–level UMI counts $Y=(Y_{1},\ldots,Y_{p})\in\mathbb{R}^{p}$, gRNA binding indicators $D=(D_{1},\ldots,D_{p})\in\{0,1\}^{p}$, and technical confounders $X=(X_{1},\ldots,X_{J})\in\mathbb{R}^{J}$, all of which are observable. Each cell receives a perturbation on a single gene. ARGEN relies on a classical measurement model (sarkar2021separating) that generates UMI counts $(Y_{1},\ldots,Y_{p})$ from latent true expression levels $(\mu_{1},\ldots,\mu_{p})$. Each $\mu_{j}$ directly receives the perturbation treatment and is influenced by both observed ($X$) and unobserved confounders $(U=(U_{1},\ldots,U_{m}))$. The gRNA binding indicator is also associated with the observed technical factors $X$. b. A simplified DAG with $p=4$, representing the structural equation model of ARGEN. The goal of ARGEN is to learn the (red) edges among the unobserved true expression variables. For simplicity, $X$ is omitted from the visualization. c. Overview of ARGEN’s DAG search procedure. Left: ARGEN first estimates the descendant sets of each $\mu_{j}$ by testing (4) across $k\neq j$. The estimated $\{\mathrm{des}(j)\mid j\in[p]\}$ are then passed to Algorithm 1. Right: ARGEN subsequently identifies the parent set $pa(j)$ for each $j\in[p]$, proceeding from the bottom layer to the top within Algorithm 1. Parents are estimated via the following steps: (i) construct proxy measurements $\eta_{k}$ for the true expression $\mu_{k}$ for all $k\in anc(j)$; (ii) solve a QMLE to estimate $\theta_{jk}$ for each gene $j$; and (iii) adjust the p-values of $\theta_{jk}$, denoted by $p_{\theta_{jk}}$, using Online FDR (zrnic2020power) across $j\in[p]$ to obtain the final estimate of $pa(j)$.
@@ 2.1 Data observed in perturb-seq experiments and DAG
We present schematic overview of the data and our proposed method in Figure 1. We denote the unique molecular identifier (UMI) count of $p$ genes of a cell in a Perturb-seq data as $Y_{1},\cdots,Y_{p}$. We also denote guide RNA (gRNA) perturbation indicator $D_{j}\in\{0,1\}$, where $j\in\{1\cdots,p\}$, and $D_{j}=1$ indicating that the cell is induced by a guide that perturbs the gene $Y_{j}$ $\forall j\in[p]$. We further assume $\sum_{j=1}^{p}D_{j}\in\{0,1\}$, meaning that we target a single gene at a time or do not target any genes. A cell with $D_{1}=D_{2}\cdots=D_{p}=0$ indicates that the cell is induced by a non-targeting guide that perturbs no genes. We define the unobserved biological confounders such as accessibility level of enhancer regions or even unmeasured expression level of other genes as $U=(U_{1},\cdots,U_{m})^{T}$ that influences gene expression levels. When the Perturb-seq experiments are conducted on multiple cell types that are not pre-labeled by the data generators, the cell type label can also function as unmeasured confounders. Lastly, we define the observable cell-specific covariates as $X=(X_{1},\cdots,X_{J})^{T}$, including technical factors such as batch or GEM (short for Gel bead-in-EMulsion) group labels, which influence both the expression level and $D_{j}$ (Fig. 1a).
We use DAG to represent the causal network among the expression levels of $p$ genes. Specifically, a DAG $G$ is a pair $(V,E)$ consisting of a finite set of nodes $V(=V(G))=\{1,\ldots,p\}$ and a set of directed edges $E\subseteq V\times V$ such that the graph contains no directed cycles. That is, there does not exist a sequence of distinct nodes $i_{1},i_{2},\ldots,i_{k}$ with directed edges $i_{1}\rightarrow i_{2}\rightarrow\cdots\rightarrow i_{k}\rightarrow i_{1}.$ Each node $j\in V$ typically represents a random variable $\mu_{j}$, and a directed edge $i\rightarrow j$ encodes a potential direct causal or regulatory influence of $\mu_{i}$ on $\mu_{j}$. The acyclicity property implies that the variables can be topologically ordered so that every variable depends only on its predecessors in that order. This property allows one to represent the joint distribution of $\mu_{j}$s via a structural equation model (SEM) of the form
@@ 2.2 Classical measurement model for scRNA-seq and its extension to Perturb-seq data
We consider the standard and commonly applied measurement model for scRNA-seq datasets (huang2018saver, townes2019feature, sarkar2021separating). Specifically, we define the hidden expression level of a gene $j\in[p]$ in a cell $i\in[N]$ as $\lambda_{ij}$ and consider the observed UMI count of the gene in the cell as a random Poisson draw from the distribution
@@
where $\ell_{i}$ denotes the sequencing depth of a cell $i$, which can be typically estimated by the total UMI count for a cell $i$ (vallejos2017normalizing) or by alternative approaches such as scran (l2016pooling). We can consider different types of (prior) expression models for $\lambda_{ij}$ (sarkar2021separating), such as
@@
where $\epsilon_{ij}$ is a $i.i.d$ mean zero noise independent of other variables. In addition, our target parameter $\theta_{jk}$ quantifies the influence of gene $k$ to $j$ and $\tau_{j}$ quantifies how effective the perturbation $D_{ij}$ is for gene $j$. We also note that the binary indicator $D_{ij}$ can depend on the observed technical factors $X_{i}$ or $l_{i}$ (barry2024robust), which we will specify in the later sections. On the other hand, we assume independency between $D_{ij}$ and $U_{i}$ given $X_{i}$ and $l_{i}$, which implies that the hidden biological factors have less to do with whether or not a gene is perturbed. Lastly, notice that the unobserved confounder $U_{i}$ can influence $\log{\mu_{ik}}$ for $k\in pa(j)$ and therefore $\theta_{jk}$ is not in general identifiable, unless $\gamma_{j}=0$. Throughout, we suppress the cell index $i$ when no ambiguity arises.
Note that the model provides a super-DAG involving $\mu_{j},Y_{j},X_{j},U_{j},D_{j}$ as participating nodes (Figure. 1b). Here, we focus on learning the DAG $G$ among the true expressions $\mu_{j}$s (red subgraph in Figure. 1b), and our goal is to correctly estimate $\{pa(j)|j\in[p]\}$ and $G$.
@@ 3 Identifiability of the Causal Network
Before estimating the DAG among the genes, an DAG identification result must be established under the assumed mdoel and settings. We provide detailed identification results in Section LABEL:supp-sec:DAG_identification in Supplemental Materials with several theoretical arguments. One key point is that we need the following much weaker assumption than the classical faithfulness (pearl2009causality, uhler2013geometry) or influentiality (tian2013causal, yang2018characterizing), which turns out to be sufficient for recovering the ancestral information of the entire DAG.
@@
We show in Theorem LABEL:supp-thm:des that the ancestor nodes of $j$, denoted as $anc(j)$ and the descendant nodes, noted as $des(j)$ for $\forall j\in[p]$, can be identified. This is justified by introducing a new concept called exlusive directed path (Definition LABEL:supp-def:exclusive_directed_path), which denotes a single unique directed path between two nodes. We then show the existence of exclusive directed path in between any two nodes in Lemma LABEL:supp-lem:exclusive_path and that influentiality can be achieved in the exclusive directed path in Lemma LABEL:supp-lem:inf_exclusive_path. This leads to the following key identifiability results.

### v2

# Confounder-robust causal discovery and inference in Perturb-seq using proxy and instrumental variables

@@ Abstract
Emerging single-cell technologies that integrate CRISPR-based genetic perturbations with single-cell RNA sequencing, such as Perturb-seq, have substantially advanced our understanding of gene regulation and causal influence of genes. While Perturb-seq data provide valuable causal insights into gene-gene interactions, statistical concerns remain regarding unobserved confounders that may bias inference. These latent factors may arise not only from intrinsic molecular features of regulatory elements encoded in Perturb-seq experiments, but also from unobserved genes arising from cost-constrained experimental designs. Although methods for analyzing large-scale Perturb-seq data are rapidly maturing, approaches that explicitly account for such unobserved confounders in learning the causal gene networks are still lacking. Here, we propose a novel method to recover causal gene networks from Perturb-seq experiments with robustness to arbitrarily omitted confounders. Our framework leverages proxy and instrumental variable strategies to exploit the rich information embedded in perturbations, enabling unbiased estimation of the underlying directed acyclic graph (DAG) of gene expressions. Simulation studies and analyses of CRISPR interference experiments of K562 cells demonstrate that our method outperforms baseline approaches that ignore unmeasured confounding, yielding more accurate and biologically relevant recovery of the true gene causal DAGs.
@@ 1 Introduction
A central objective in modern genomics is to recover the directed structure of gene–gene regulation that governs cellular function. Single-cell CRISPR perturbation assays coupled with scRNA-seq, popularized by Perturb-seq, deliver interventional readouts at single-cell resolution across thousands of targeted knockdowns or activations in pooled format (Dixit et al. 2016, Replogle et al. 2022). Beyond revealing heterogeneous and context-specific responses at the single cell levels, these technologies create large-scale datasets that, in principle, permit causal discovery of causal gene networks (CGNs) rather than mere association maps. Such CGNs are often presented as directed acyclic graphs (DAGs), where nodes are genes and directed edges present regulatory relationships.
@@
From a statistical perspective, learning DAGs from observational or interventional data has been studied widely in literature. Classical constraint-based procedures such as the PC algorithm (Spirtes et al. 2000) rely on conditional independence tests to identify Markov equivalence classes under assumptions of causal sufficiency and faithfulness. Score-based approaches such as the Greedy Equivalence Search (GES) (Chickering 2002) optimize penalized likelihoods over equivalence classes, while hybrid approaches combine constraint and score elements. Continuous optimization frameworks such as NOTEARS (Zheng et al. 2018) and its nonlinear extensions (Zheng et al. 2020) recast DAG learning as smooth optimization problems, enabling scalable inference in high dimensions. Other methods, including those based on additive noise models (Peters et al. 2014), provide identifiability for non-Gaussian or nonlinear settings. While these frameworks have advanced causal discovery in observational data, they generally assume that all confounders are observed and that interventions are perfectly applied and measured. When interventional data are available, causal discovery can, in principle, be strenghtened. Algorithms such as GIES (Hauser and Bühlmann 2012) extend GES to integrate intervention targets for orienting edges and improving identifiability, and active-learning variants optimize which interventions to perform next (Agrawal et al. 2019).
@@
However, these methods typically presume that all participating nodes of the underlying DAG are observed in the data, without omitted factors that can alter the distributions of confounded nodes. In pooled CRISPR single-cell screens, this assumption is likely to be violated. In particular, biological heterogeneity, including cell cycle phase, chromatin accessibility, and lineage differences, introduces additional latent variation that is correlated with the transcriptomic profiles of cells. Moreover, due to cost considerations, researchers implementing CRISPR experiments are often interested in intervening on a small set of key genes, which can produce incomplete datasets with omitted genes. Ignoring unobserved factors may result in biased inference of causal relationships in DAGs. In addition, technical covariates such as sequencing depth, capture efficiency, and batch effects influence both gene expression and gRNA detection. Moreover, the data are high-dimensional and highly unbalanced across targets, with a small number of cells per perturbation and unknown timescales for transcriptional responses. These characteristics violate assumptions underlying standard DAG-learning methods.
@@
Several computational tools have been proposed to model Perturb-seq data. The original Perturb-seq studies introduced MIMOSCA for regularized linear modeling of perturbation effects (Dixit et al. 2016). Subsequent methods such as scMAGeCK (Yang et al. 2019) extended genotype-to-multivariate-phenotype mapping for single-cell CRISPR screens. More recently, SCEPTRE (Katsevich and Roeder 2022) provided a statistically calibrated approach using conditional resampling to correct for confounding from technical covariates, substantially improving type I error control and sensitivity in detercting the affected genes from perturbation. Despite these advances, current frameworks primarily focus on estimating pairwise perturbation–response effects and do not naturally extend to reconstructing global regulatory DAGs. Moreover, robust computational methods for estimating DAGs in the presence of unobserved factors are still lacking.
@@
In this paper, we propose ARGEN for Arbitrary-confounder Robust causal GEne Network, as a statistical framework for recovering causal network among genes from Perturb-seq experiments that is robust to arbitrarily omitted confounders. ARGEN explicitly leverages the interventional design of CRISPR screens, and adopts an instrumental variable strategy that uses proxy gene expression to identify causal gene networks. Guide detections serve as instrumental variables, and proxies for hidden gene expression provide robustness against confounding. We extend the classical measurement model for scRNA-seq data (Townes et al. 2019, Sarkar and Stephens 2021) to a setting that explicitly reflects both the underlying gene regulatory network and the perturbation structure inherent to Perturb-seq data. Specifically, we employ a structural equation model (SEM) with embedded perturbation terms to represent the causal relationships among gene expressions under CRISPR interventions. Under the assumed perturbation model, we investigate conditions for identification of the ancestors and descendants of each perturbed (intervened) target gene, as well as the identification of the causal ordering and the underlying DAG. We further develop computationally efficient algorithms for estimating descendant sets and parental relationships, with provable guarantees that the resulting estimated structure satisfies the acyclicity constraint and yields a valid DAG, even in finite-sample settings.
@@ 2 Statistical Model for Perturb-seq Data
Figure 1: Schematic overview of ARGEN. a. ARGEN takes as input single-cell–level UMI counts $Y=(Y_{1},\ldots,Y_{p})\in\mathbb{R}^{p}$, gRNA binding indicators $D=(D_{1},\ldots,D_{p})\in\{0,1\}^{p}$, and technical confounders $X=(X_{1},\ldots,X_{J})\in\mathbb{R}^{J}$, all of which are observable. Each cell receives a perturbation on a single gene. ARGEN relies on a classical measurement model (Sarkar and Stephens 2021) that generates UMI counts $(Y_{1},\ldots,Y_{p})$ from latent true expression levels $(\mu_{1},\ldots,\mu_{p})$. Each $\mu_{j}$ directly receives the perturbation treatment and is influenced by both observed ($X$) and unobserved confounders $(U=(U_{1},\ldots,U_{m}))$. The gRNA binding indicator is also associated with the observed technical factors $X$. b. A simplified DAG with $p=4$, representing the structural equation model of ARGEN. The goal of ARGEN is to learn the (red) edges among the unobserved true expression variables. For simplicity, $X$ is omitted from the visualization. c. Overview of ARGEN’s DAG search procedure. Left: ARGEN first estimates the descendant sets of each $\mu_{j}$ by testing (4) across $k\neq j$. The estimated $\{\mathrm{des}(j)\mid j\in[p]\}$ are then passed to Algorithm 1. Right: ARGEN subsequently identifies the parent set $pa(j)$ for each $j\in[p]$, proceeding from the bottom layer to the top within Algorithm 1. Parents are estimated via the following steps: (i) construct proxy measurements $\eta_{k}$ for the true expression $\mu_{k}$ for all $k\in anc(j)$; (ii) solve a QMLE to estimate $\theta_{jk}$ for each gene $j$; and (iii) adjust the p-values of $\theta_{jk}$, denoted by $p_{\theta_{jk}}$, using Online FDR (Zrnic et al. 2020) across $j\in[p]$ to obtain the final estimate of $pa(j)$.
@@ 2.1 Data observed in perturb-seq experiments and DAG
We present schematic overview of the data and our proposed method in Figure 1. We denote the unique molecular identifier (UMI) count of $p$ genes of a cell in a Perturb-seq data as $Y_{1},\cdots,Y_{p}$. We also denote guide RNA (gRNA) perturbation indicator $D_{j}\in\{0,1\}$, where $j\in\{1\cdots,p\}$, and $D_{j}=1$ indicating that the cell is induced by a guide that perturbs the gene $Y_{j}$ $\forall j\in[p]$. We further assume $\sum_{j=1}^{p}D_{j}\in\{0,1\}$, meaning that we target a single gene at a time or do not target any genes. A control cell with $D_{1}=D_{2}\cdots=D_{p}=0$ indicates that the cell is induced by a non-targeting guide that perturbs no genes. We define the unobserved biological confounders such as accessibility level of enhancer regions or even unmeasured expression level of other genes as $U=(U_{1},\cdots,U_{m})^{T}$ that influences gene expression levels. When the Perturb-seq experiments are conducted on multiple cell types that are not pre-labeled by the data generators, the cell type label can also function as unmeasured confounders. Lastly, we define the observable cell-specific covariates as $X=(X_{1},\cdots,X_{J})^{T}$, including technical factors such as batch or GEM (short for Gel bead-in-EMulsion) group labels, which influence both the expression level and $D_{j}$ (Fig. 1a).
We use DAG to represent the causal network among the expression levels of $p$ genes. Specifically, a DAG $G$ is a pair $(V,E)$ consisting of a finite set of nodes $V(=V(G))=\{1,\ldots,p\}$ and a set of directed edges $E\subseteq V\times V$ such that the graph contains no directed cycles. That is, there does not exist a sequence of distinct nodes $i_{1},i_{2},\ldots,i_{k}$ with directed edges $i_{1}\rightarrow i_{2}\rightarrow\cdots\rightarrow i_{k}\rightarrow i_{1}.$ Each node $j\in V$ typically represents a random variable $\mu_{j}$ that represents the true expression of gene $j$ in a given cell, and a directed edge $i\rightarrow j$ encodes a potential direct causal or regulatory influence of $\mu_{i}$ on $\mu_{j}$. The acyclicity property implies that the variables can be topologically ordered so that every variable depends only on its predecessors in that order. This property allows one to represent the joint distribution of $\mu_{j}$s via a structural equation model (SEM) of the form
@@ 2.2 Classical measurement model for scRNA-seq and its extension to Perturb-seq data
We consider the standard and commonly applied measurement model for scRNA-seq count data (Townes et al. 2019, Sarkar and Stephens 2021). Specifically, we define the hidden expression level of a gene $j\in[p]$ in a cell $i\in[N]$ as $\lambda_{ij}$ and consider the observed UMI count of the gene in the cell as a random Poisson draw from the distribution
@@
where $\ell_{i}$ denotes the sequencing depth of a cell $i$, which can be typically estimated by the total UMI count for a cell $i$ (Vallejos et al. 2017) or by alternative approaches such as scran (Lun et al. 2016). We can consider different types of (prior) expression models for $\lambda_{ij}$ (Sarkar and Stephens 2021), such as
@@
where $\epsilon_{ij}$ is a $i.i.d$ mean zero noise independent of other variables. In addition, our target parameter $\theta_{jk}$ quantifies the influence of gene $k$ to $j$ and $\tau_{j}$ quantifies how effective the perturbation $D_{ij}$ is for gene $j$. We also note that the binary indicator $D_{ij}$ can depend on the observed technical factors $X_{i}$ or $l_{i}$ (Barry et al. 2024). On the other hand, we assume independency between $D_{ij}$ and $U_{i}$ given $X_{i}$ and $l_{i}$, i.e., $D_{ij}\perp\!\!\!\!\perp U_{i}|\ell_{i},X_{i}$, which implies that the hidden biological factors have less to do with whether or not a gene is perturbed. Lastly, notice that the unobserved confounder $U_{i}$ can influence $\log{\mu_{ik}}$ for $k\in pa(j)$ and therefore $\theta_{jk}$ is not in general identifiable, unless $\gamma_{j}=0$. Throughout, we suppress the cell index $i$ when no ambiguity arises.
The model above provides a super-DAG involving $\mu_{j},Y_{j},X_{j},U_{j},D_{j}$ as participating nodes (Figure. 1b). Here, we focus on learning the DAG $G$ among the true expressions $\mu_{j}$s (red subgraph in Figure. 1b), and our goal is to correctly estimate $\{pa(j)|j\in[p]\}$ and $G$.
@@ 3 Identifiability of the Causal Network
Before estimating the DAG among the genes, a DAG identification result must be established under the assumed mdoel and settings. We provide detailed identification results in Section LABEL:supp-sec:DAG_identification in Supplemental Materials with several theoretical arguments. One key point is that we need the following much weaker assumption than the classical faithfulness (Pearl 2009, Uhler et al. 2013) or influentiality (Tian and Pearl 2013, Yang et al. 2018), which turns out to be sufficient for recovering the ancestral information of the entire DAG.
@@
We show in Theorem LABEL:supp-thm:des that the ancestor nodes of $j$, denoted as $anc(j)$ and the descendant nodes, noted as $des(j)$ for $\forall j\in[p]$, can be identified. This is justified by introducing a new concept called exlusive directed path (Definition LABEL:supp-def:exclusive_directed_path), which denotes a single unique directed path between two nodes. We then show the existence of such path in the subgraph between any two nodes in Lemma LABEL:supp-lem:exclusive_path and that influentiality can be achieved in the exclusive directed path in Lemma LABEL:supp-lem:inf_exclusive_path. This leads to the following key identifiability results.